OpenStudy (anonymous):

which of the following is the simplified form of (x^2)-16/x-4 1) x+4, with the restriction x≠ -4 2) x-4, with the restriction x≠ -4 3) x+4, with the restriction x≠ 4 4) x-4, with the restriction x≠ 4

OpenStudy (anonymous):

\[x ^{2}-16/(x-4) = (x-4)(x+4)/(x-4) = x+4\]

OpenStudy (anonymous):

would that be 1 for 3?

OpenStudy (anonymous):

3